What is the speed of an electron with kinetic energy 760 eV ?

What is the speed of an electron with kinetic energy 0.52 keV ?
Answer #1

Kinetic Energy = 1/2 mv2 = eV

Speed, v = V2eV/m)

Mass of electron, m = 9.11*10-31 kg

Charge on electron, e = 1.6*10-19 C

760 eV = 760 * 1.6 * 10-19 = 1.216 * 10-16 J

0.52 keV = 520 * 1.6 * 10-19 =  8.32 * 10-17 J

v1 = (2* 1.216 1016)/(9.11 10-31 = 1.6339 * 107 m/s

v2 = (2* 8.32 10-17)/(9.11 10-31 = 1.3515 * 107 m/s